# Understanding the Application of Utility Theory in Robotics and Artificial Intelligence: A Survey

### v1

@@ II-A Utility Theory
Utility theory is postulated in economics to explain the behavior of individuals based on the premise that individuals can consistently rank the order of their choices depending on their needs or preferences. This theoretical approach aims to quantify an agent’s degree of preference across a set of available alternatives and understand how these preferences change when an agent faces uncertainty about which alternative it will receive [8].
@@
To describe the interactions between multiple utility-theoretic agents, we use the specific utility function to analyze their preferences and rational action. The utility function is a mapping from states of the world to real numbers, which are interpreted as measures of an agent’s level of happiness (needs) in the given states. If the agent is uncertain about its current state, the utility is defined as the expected value (Eq. (1)) of its utility function for the appropriate probability distribution over the states [8]. From the perspective of the connection between a decision-maker and its preference, a decision-maker would rather implement a more preferred alternative (act, course of action, strategy) than one that is less preferred [9].
@@
Generally, utility serves as a unified concept in value systems to describe the capacity of action, behavior, and strategy to satisfy the agents’ desires or needs in their decision-making. Furthermore, modern utility theory has its origins in the theory of expected value, and the value of any course of action could be determined by multiplying the gain realized from that action by the likelihood of receiving that gain [10].
II-B Basic Concepts in Agent’s Motivation
@@
Supposing AI agents can sense various environments, monitor their behaviors, and get internal and external information through perception $P(t)$ in the given time. And the data can generally be categorized into two classes: different variables in the environments (like light, sound, color, etc.) and agents’ physical structure and operation (such as arm extension, contact in leg, etc), the operation of the cognitive processes within the robot (learning error, perceptual novelty, etc.) [11]. Then, the set of values transmits to the agent’s actuators making up the actuation vector $A(t)$ in a given time. Furthermore, we can define a series of specific concepts as follow:
@@
Especially, [14] presents a model for a generic intrinsically motivated agent shown in Fig. 3  [^2] Here, intrinsic motivation components modulate behavioral components to connect perception components (sensors) and actuators.. Through sensors generating an environment state at each time step, the agent will form various goals and corresponding expected utilities in motivation components. Then, the behavioral components will make decisions based on specific goals to modify the agent’s internal state and generate diverse strategies and actions to achieve goals by triggering actuators to interact with the environment changing the external state.
II-C Needs-driven Behaviors
@@
In nature, from cell to human, all intelligent agents represent different kinds of hierarchical needs such as the low-level physiological needs (food and water) in microbe and animal; the high-level needs self-actualization (creative activities) in human being [15]. Different levels of needs stimulate all agents to adopt diverse behaviors achieving certain goals or purposes and satisfying their various needs in the natural world [16]. They also can be regarded as self-organizing behaviors [17], which keep on interacting with the environments and exchanging information and energy to support the system’s survival and development.
@@
Specifically, the needs describe the necessities for a self-organizing system to survive and evolve, which arouses an agent to action toward a goal, giving purpose and direction to behavior and strategy. The dominant approach to modeling an agent’s interests or needs is Utility Theory. [18] introduced the self-sufficient robots’ decision-making based on a basic work cycle – find fuel – refuel. They utilized some utility criterion needs, such as utility maximization, to guide robots performing opportunistic behaviors [19].
@@
From the utility theory perspective, needs are regarded as innate values driving agents to interact with environments. Especially the ultimate payoff for animals is to maximize inclusive fitness utilizing some evolutionary strategy, and for humans is to maximize profitability through some marketing strategy [20, 21]. If we consider the ecological status of AI agents, like robots, accomplishing tasks in the real world, their situation is similar to our ecological situation [22].
Especially reinforcement learning (RL) is an excellent model to describe the needs-driven behaviors of AI agents. The essence of RL is learning from interaction based on reward-driven (such as utilities and needs) behaviors, much like natural agents. When an RL agent interacts with the environment, it can observe the consequence of its actions and learn to change its behaviors based on the corresponding rewards received. The theoretical foundation of RL is the paradigm of trial-and-error learning rooted in behaviorist psychology [23]. In RL, the policy dictates the actions that the agent takes as a function of its state and the environment, and the goal of the agent is to learn a policy maximizing the expected cumulative rewards/utilities in the process. Fig. 1 illustrates the RL process of rewards-driven behaviors [24].
II-D Utility-orient Needs Paradigm Systems
Fig. 2 represents the relationships among single-agent, multi-agent, and humans based on the paradigm of the utility-orient needs. More specially, before an individual agent cooperates with group members, it needs to fit some basic needs, such as having enough energy, guaranteeing other agents’ safety, etc. Then it requires corresponding capabilities to collaborate with other agents, satisfy task requirements, and adapt to dynamically changing environments (Fig. 2 shows four categories of robot, carrier, supplier, observer, and executor, working as a team in the middle graph.). Especially trust among agents can be evaluated based on their current status and needs, performance, knowledge, experience, motivations, etc.
Furthermore, when AI agents work with humans, they should satisfy humans’ needs and assist humans in fulfilling their missions efficiently. Therefore, it is necessary to build a unified utility paradigm describing the needs of AI agents and humans in the common ground for evaluating their relationships like trust, which is also the pre-condition of safety, reliability, stability, and sustainability in cooperation. In the following sections, we will review existing literature from the perspective of single-agent systems, multi-agent systems (MAS), trust among agents, and human-robot interaction (HRI) to support the proposed utility-orient needs paradigm systems.
III Single-Agent Systems

### v3

@@ 2.3.1 Basic Concepts
The dominant approach to modeling an agent’s interests or needs is utility theory. This theoretical approach aims to quantify an agent’s degree of preference across a set of available alternatives and understand how these preferences change when an agent faces uncertainty about which alternative he will receive [234].
@@
To describe the interactions between multiple utility-theoretic agents, we use the specific utility function to analyze their preference and rational action. The utility function is a mapping from states of the world to real numbers, which are interpreted as measures of an agent’s level of happiness (needs) in the given states. If the agent is uncertain about its current state, the utility is defined as the expected value of its utility function for the appropriate probability distribution over states [234]. From the perspective of the connection between a decision maker and its preference, a decision maker would rather implement a more preferred alternative (act, course of action, strategy) than one that is less preferred [83].
Furthermore, expected utility theory states that the decision maker choose between risky or uncertain prospects by comparing their expected utility values, that is, the weighted sums obtained by adding the utility values of outcomes multiplied by their respective probabilities [180]. There are two kinds of theories based on the standard distinction between risk and uncertainty: Subjective expected utility theory (SEUT) and Von Neumann–Morgenstern theory (VNMT).
Figure 8: Illustration of the value systems of human based on neuromodulators [106] and the AI robot’s value systems from utility theory perspective.
Subjective expected utility theory:
The theory of subjective expected utility describes the attractiveness of an economic opportunity as perceived by a decision-maker in the presence of risk, which combines two subjective concepts: a personal utility function and a personal probability distribution (usually based on Bayesian probability utility theory) [227, 118].
In SEUT, if the decision maker adheres to axioms of rationality, believing an uncertain event has possible outcomes $\{x_{i}\}$ each with utility $u(x_{i})$, then the person’s choices can be explained as arising from this utility function combined with the subjective belief that there is a probability of each outcome, $P(x_{i})$. And the subjective expected utility is Eq. (1).
@@
The decision relies on the higher subjective expected utility, which means agents may make different decisions because of their various utility functions or beliefs about the probabilities of different outcomes. Moreover, [181] further discussed the connection between ethical and utility theory describing the complex social behaviors in human.
Von Neumann–Morgenstern utility theory:
In decision theory, the Von Neumann–Morgenstern utility theory (VNMT) is also known as the expected utility hypothesis, which shows that, under certain axioms [^3] The four axioms of VNM-rationality are completeness, transitivity, continuity, and independence. [259] of rational behavior, a decision-maker faced with risky (probabilistic) outcomes of different choices will behave as if the agent is maximizing the expected value of some function defined over the potential outcomes at some specified point in the future [259].
Maslow’s hierarchy of human needs:
From the psychological perspective, Maslow’s hierarchy of needs [163] is used to study how humans intrinsically partake in behavioral motivation. It is often portrayed in the shape of a pyramid, with the largest, most fundamental needs at the bottom, and the need for self-actualization and transcendence at the top. In other words, the idea is that individuals’ most basic needs must be met before they become motivated to achieve higher-level needs. Maslow used the terms physiological, safety, belonging and love, social needs or esteem, and self-actualization to describe the pattern through which human motivations generally move and the conditional dependent relationships between each level’s needs. In other words, if the agent’s motivation wants to arise at the next stage, the current level’s needs (utilities) must satisfy a certain amount. Furthermore, for more advanced intelligent agents, such as human beings, many motivations from different levels of needs can occur simultaneously. Instead of focusing on a certain need at any time, [164] stated that a specific need might dominate the decision at a certain time, and the likelihood of the dominant need will follow the order presented in the needs hierarchy.
Generally speaking, if we can build a model describing the agent’s hierarchy of needs based on the expected utility theory, we can develop a more intelligent single agent and MAS representing more complex behaviors to adapt to various dynamic and complex environments.
2.3.2 Experienced Utility in Cognitive Process
People make decisions to gain pleasure and avoid pain [28]. In the view of neoclassical economics, people try to maximize utility to fit their individual preferences or needs, which starts with a hedonic approach to utility that aims to produce the most pleasure and the least pain [30]. This view can also be explained in terms of neuroscience. Neuromodulators in the central nervous system include serotonin, acetylcholine, dopamine, and norepinephrine, which transmit signals across multiple neurons. They build our reward mechanism (Fig. 8) to induce subjective feelings of pleasure and contribute to positive emotions [231]. For an individual, a good decision is to maximize the hedonic or pleasurable experience of the chosen outcome  [^4] Outcomes that generate a pleasure impact elicit a constellation of objective responses (including affective behavioral reactions, physiological autonomic, and brain limbic reactions) as well as in humans at least, subjective feelings reported as pleasure., termed Experienced Utility [82]. Experienced utility or subjective pleasure involves cognitive construal of future goals [115]. The process generates our motivation, needs, and beliefs [73], described as a cognitive form of anticipated utility. From a neuroeconomic perspective, decision-making based on experienced utility is associated with widespread neural activation across a diffuse circuit network involving many brain structures, which helps us better understand the relationships between cognition and the brain [96].
2.3.3 Utility Value Systems of Cognition
A value system facilitates the capacity of a biological brain to increase the likelihood of neural responses to an external phenomenon [26]. Value presents a subjective point of view to measure the effort of an agent willing to expend to obtain rewards or to avoid punishment from goals. Due to its biological inspiration, it is termed innate, inherent, or motivational value [173], which reflects a subjective evaluation of the sensory space. In the life of an agent, the agent starts from inherent values, and its acquired values are modified through interaction with various environments and knowledge accumulation in learning. So the acquired value is thus activity-dependent and allows the value system to become sensitive to stimuli that are not able to trigger a value-related response by themselves [240]. [238] explains that a vertebrate’s value system is realised in part through the roles of neuromodulators and a form of brain plasticity called long-term potentiation, which develops acquired value systems based on the reward mechanism formed through neuromodulators, as we discussed above (Fig. 8). To evaluate an agent’s innate values and motivations, utility theory aims to quantify an agent’s degree of preference across a set of available alternatives and understand how these preferences change when the agent faces uncertainty about which alternative it will receive [234, 96]. Furthermore, modern utility theory has its origins in the theory of expected value, and the value of any course of action can be determined by multiplying the gain from that action by its likelihood [97].
Specifically, neurons process utility rewards derived from environmental information via distinct brain structures. They detect the expectation of imminent rewards or anticipated utility in the value and identity of encountered stimuli, as reflected in the control of goal-directed behaviour. The stimuli will be labelled with a positive value proportional to the degree of the reward signal. This information helps construct neuronal representations that enable subjects to anticipate future utility rewards based on prior experience. Through cognitive appraisals of the experience utility, the individual can perform subjective cost-benefit calculations to determine the level of effort required and execute effortful behavioural responses. To balance the contingencies and utility rewards, it naturally adopts adaptive behaviour to respond to the changing environments. This process will gradually form its unique value direction and preferences, represented as a distribution of needs [279], termed value systems [231].
@@ 2.3.3 Utility Value Systems of Cognition
To summary, utility serves as a unified concept in value systems to describe the capacity of action, behavior, and strategy to satisfy the agents’ desires or needs in their decision-making.
2.4 Utility Theory based Cognitive Modeling in Robotics
2.4.1 Value Systems of Cognitive Robots
Figure 9: The illustration of the value system in the cognitive architecture of SAIL robot [111].
Table 2: Aspects of the Value System Design for Cognitive Robots
References Robot Value Type1 Data Structure2 Initial Strategy3 Update Strategy4
[238] NOMAD Extrinsic: Taste Neural Network Experience Synaptic Rules
[209] Kephera Extrinsic: Fear & Hunger Feedforward Networks Predefined Euler method
[63] CARL-1 Intrinsic: Explore Simulated Neurons Random Weight Rules
[34] Simulation Intrinsic: Novelty Neural Units Predefined Multiple
[175] Lego Mindstorms NXT Extrinsic: Behavior Cycles Neural Network Random Q-Learning
[161] Simulation Extrinsic: Empowerment Neural Network Random Information Maximisation
[111] SAIL Intrinsic: Novelty Regression Tree Random Q-Learning
[71] Cyber Rodent Extrinsic: Reproduction Various Various Reinforcement Learning (RL)
[85] iCub Humanoid Intrinsic: Curiosity Markov decision process (MDP) MoBeE [86] RL
[139] Robot Arm Intrinsic: Novelty Arrays Mapping
[128] uBot-6 Intrinsic: Surprise Aspect Transition Graph (ATG) Initial ATG Provided Recursive Bayesian Estimation
[226] Simulation Extrinsic: Changing Environment Multiple Multiple Multiple
[80] Simulation Extrinsic: Changing Light Leaky Integrator Units Experience Hebbian Learning
[101] Lego Mindstorms Arm Extrinsic: Predicting Error Neural Network Random iNAC5
[195] Sony AIBO dog Intrinsic: Competence Progress Exemplars Random Nearest Neighbour Algorithm
[188] Katana Robot Arm Intrinsic: Learning Progress Parametric Representation Predefined LSPI6
[64] Kinect Intrinsic: Novelty Random forest Random Decision Tree
[281] [282] Simulation Combined: Energy & Saftey & Tasks Needs Hierarchy Predefined Behavior trees
[284] Simulation Combined: Energy & Victim Needs Hierarchy Predefined Relative Entropy
[285] Robotarium [205] Combined: Survival & Pursuit Needs Hierarchy Predefined GUT7
[279] Simulation Combined: Survival & Tasks Needs Hierarchy & Neural Network Predefined Innate-Values-Driven RL (IVRL)
  • 1
    Value Type: Intrinsic Motivations, Extrinsic Motivations, and Combined Motivations.
  • 2
    Data Structure: The type of value network is represented in a value system.
  • 3
    Initial Strategy: The initial selected values affect early behaviors and innate values.
  • 4
    Update Strategy: The mathematical process updates the acquired values.
  • 5
    iNAC: Incremental atural actor-critic [32]
  • 6
    LSPI: Least squares policy iteration [133]
  • 7
    GUT: Game-theoretic utility tree [286]
The value measures an agent’s effort to obtain a reward or avoid punishment. It is not hard-wired for an AI agent (like robots), even a biological entity, and the specific value system achieved through experience reflects an agent’s subjective evaluation of the sensory space. And the value mechanisms usually have been defined as the expected values, particularly in uncertain environments. Moreover, the innate value reflects an agent’s subjective evaluation of the sensory space, but the acquired value is shaped through experience during its development. A value system plays a crucial role in developing human-like intelligence in robotic systems, reflecting a robot’s self-awareness, trust, ethics, knowledge, motivations, and needs. [26] defines the artificial value system for the autonomous robot as a mechanism that can reflect the effects of robots’ experience on future behaviors, such as perception, planning, exploration, and learning. Fig. 9 illustrates that [111] introduced the modulating effect of the value system to guide the behavior of an autonomous system in the SAIL robot cognitive architecture. Moreover, the robotic value system demonstrates its capacity to plan action upon detecting salient stimuli, performing action planning after analyzing its internal and external context [112, 237]. Here, the robotic internal context describes its knowledge developed through past experiences, and the external context expresses its preference for a planned action in the current environment.
Existing computational value systems exhibit various distinctions between value and reward, which we can abstract as a network of nodes and edges in the fig. 10. Specifically, the input nodes receive information from the external environment, and the output nodes will trigger a behavioural response by the robot. The edge between each pair of nodes describes the strength of the connection, represented by a scalar weight. The path between the input and output nodes corresponds to the total strength through the value appraisal, which indicates the likelihood of a certain behavioural reaction to an external stimulus detected by input nodes. Therefore, a value system can be regarded as a sparse or dense network, presented as generic graph-like structures, such as a tree or tabular structure. Moreover, the Internal nodes can be neurons, decision nodes, observations, contexts, clusters, or parameters. Through updating different formulae, the system can adjust the reinforcement or decay of weighted edge connections.
Many computational models involve the term “value” [89], such as innate values in the agent’s initial state and acquired values throughout the agent’s life [239]. Considering primates’ value systems as a benchmark, [231, 74] summarize the properties of an artificial value system for cognitive robots as follows:
Figure 10: The illustration of the value system as a network of nodes and edges [173].
  • Prediction: According to external information, the value system can apply reasoning and knowledge to analyze the consequences of an action before executing it. It means that the value network (Fig. 10(a)) should have paths to link sensory stimuli, behavioral responses, and expected subsequent sensory stimuli (Fig. 10(b)).
  • Task non-specific: An artificial value system should guide a robot to learn in a general and task-agnostic manner without human intervention. In the early stages, robots can receive limited support in developing their value systems. However, the systems need to have the capability to self-evolve in response to interactive environments, much as humans do [^5] Instructional scaffolding is a widely used mechanism in primates (especially humans) to boost value system development during the very early stage, but the supervision provided through scaffolding is expected to decay gradually to facilitate natural cognitive development [260]..
  • Self-regulation and Development: Self-regulation explains how people consciously manage their thoughts, emotions, and behaviors to achieve personal goals, which involves emotional regulation, behavioral control, and cognitive management [119]. It helps us develop reasonable value systems that adapt to current environments. Like a biological entity, the artificial agent also needs a self-regulation mechanism, including both innate and acquired components, to develop suitable value systems for sustainable development. Innate value is usually modeled as a prior over utility values, assigning different weights to actions. Correspondingly, acquired value will be generated over the long term throughout the interactive process. An artificial value system must have the plasticity to continuously encode feedback from its interactions with the environment, and the experience of this encoding should facilitate the emergence of adaptive behavior.
  • Value-based learning: Learning plays a critical role throughout the development of artificial value systems, which should be able to generate modulating signals to adapt various learning parameters, such as bias (positive or negative) and learning rate [26].
Table. 2 lists design aspects and summarizes the corresponding research on value system design for autonomous robots. However, studies on designing artificial value systems with the goal of developing human-like intelligence in robotic systems remain very limited, especially in modeling human-like motivation mechanism in robots, quantifying value feedback from environments related to robotic current status and further development, and building self-upgrade artificial value systems in robots to adapt to dynamic changing environments. In the following parts, we provide the current methology about motivation modeling in robots from reviewed papers.
2.4.2 Basic Concepts in Agent’s Motivation
From the organisms’ perspective, [202] defines the motivations of an intelligent agent as having three important functions: Selection drives the system to choose a suitable behavior to achieve the most important current needs/goals; Energy provides enough power to execute the selected behavior; and Learning generates learning signals to change behavior. Furthermore, [21, 22] proposed a more operational definition: regarding functions, extrinsic motivations have the overall function of driving behavior and learning for the acquisition of material resources, like biological needs: hunger and thirst. In contrast, intrinsic motivations are processes that can drive the acquisition of knowledge and skills in the absence of extrinsic motivations, such as curiosity, surprise, novelty, and success at accomplishing the agent’s own goals. In cognitive robotics, researchers define intrinsic motivations as a means of reward to influence the accumulation of value. Different approaches include using only intrinsic rewards or combining intrinsic and extrinsic rewards in various studies.
@@ 2.4.2 Basic Concepts in Agent’s Motivation
However, the traditional AI agent lacks specific goal-setting to drive its motivations. Therefore, an intelligent cognitive architecture needs to involve the motivational system, integrating a series of mechanisms to support the agent’s self-discovery and self-selection of goals based on the drive vector established by evolution or by the system designer. Supposing AI agents can sense various environments, monitor their behaviors, and get internal and external information through perception $P(t)$ in the given time. And the data can generally be categorized into two classes: different variables in the environments (like light, sound, color, etc.) and agents’ physical structure and operation (such as arm extension, contact in leg, etc), the operation of the cognitive processes within the robot (learning error, perceptual novelty, etc.) [220]. Then, the set of values transmits to the agent’s actuators making up the actuation vector $A(t)$ in a given time. Furthermore, we can define these terms to help frame and understand related concepts in the cognitive modeling of motivation for the AI agent.
@@
[176] presents a model of a generic intrinsically motivated agent, in which intrinsic motivation components modulate behavioral components to connect perception components (sensors) and actuators, as shown in Fig. 11. Through sensors generating an environment state at each time step, the agent will form various goals and corresponding expected utilities in motivation components. Then, the behavioral components will make decisions based on specific goals to modify the agent’s internal state and generate diverse strategies and actions to achieve goals by triggering actuators to interact with the environment changing the external state.
2.4.3 Needs-driven Behaviors
@@
In nature, from cell to human, all intelligent agents represent different kinds of hierarchical needs such as the low-level physiological needs (food and water) in microbes and animals; the high-level needs for self-actualization (creative activities) in human beings [163]. Different levels of needs stimulate all agents to adopt diverse behaviors achieving certain goals or purposes and satisfying their various needs in the natural world [289]. They also can be regarded as self-organizing behaviors [24], which keep on interacting with the environments and exchanging information and energy to support the system’s survival and development.
@@
Specifically, the needs describe the necessities for a self-organizing system to survive and evolve, which arouses an agent to action toward a goal, giving purpose and direction to behavior and strategy. For example, [170] introduced the self-sufficient robots’ decision-making based on a basic work cycle – find fuel – refuel. They utilized some utility criterion needs, such as utility maximization, to guide robots performing opportunistic behaviors [172].
@@
From the utility theory perspective, needs are regarded as innate values or motivations that drive agents to interact with their environments, as we discussed previously. Especially the ultimate payoff for animals is to maximize inclusive fitness utilizing some evolutionary strategy, and for humans is to maximize profitability through some marketing strategy [177, 171]. If we consider the ecological status of AI agents, like robots, accomplishing tasks in the real world, their situation is similar to our ecological situation [169].
2.4.4 Utility-orient Needs Paradigm Systems
For the above concepts to be cognitively applicable to the intelligent system, their biology-inspired needs are somehow associated with its actuation, which drives it to fulfill its motivations or satisfy its drivers [219]. Moreover, before an individual agent cooperates with group members, it needs to fit some basic needs, such as having enough energy, guaranteeing other agents’ safety, etc. Then it requires corresponding capabilities to collaborate with other agents, satisfy task requirements, and adapt to dynamically changing environments [282]. Therefore, it is necessary to build a unified utility paradigm describing the needs of AI agents and humans in the common ground for evaluating their relationships like trust, which is also the pre-condition of safety, reliability, stability, and sustainability in cooperation [283].
[…]
Figure 13: The illustration of the various aspects of the value system development [14] from a biological brain [96] to an artificial brain.
Similar to Maslow’s hierarchy of needs, which describes humans’ motivations and value systems, we can also organize autonomous robots’ motivations as a hierarchical structure, represented as a conditional dependence network [284, 286], such as a Bayesian network. The lowest (first) level is the safety features of the agent (e.g., features such as collision detection, fault detection, etc., that assure safety to the agent, human, and other friendly agents in the environment). The safety needs (Eq. (4)) can be calculated through its safety feature’s value and corresponding safety feature’s probability based on the current state of the agent. After satisfying safety needs, the agent considers its basic needs (Eq. (5)), which includes features such as energy levels, data communication levels that help maintain the basic operations of that agent. Only after fitting the safety and basic needs, an agent can consider its capability needs (Eq. (6)), which are composed of features such as its health level, computing (e.g., storage, performance), physical functionalities (e.g., resources, manipulation), etc.
At the next higher level, the agent can identify its teaming needs (Eq. (7)) that accounts the contributions of this agent to its team through several factors (e.g., heterogeneity, trust, actions) that the team needs so that they can form a reliable and robust team to successfully perform a given mission.
Ultimately, at the highest level, the agent learns some skills/features to improve its capabilities and performance in achieving a given task, such as Reinforcement Learning. The policy features (Q table or reward function) are accounted into its learning needs expectation (Eq. (8)). The expectation of agent needs at each level are given below:
$\begin{split}Safety~~Needs:N_{s_{j}}=\sum_{i=1}^{s_{j}}S_{i}\cdot\mathbb{P}(S_{i}|X_{j},T);\end{split}$ (4)
$\begin{split}Basic~~Needs:N_{b_{j}}=\sum_{i=1}^{b_{j}}B_{i}\cdot\mathbb{P}(B_{i}|X_{j},T,N_{s_{j}});\end{split}$ (5)
$\begin{split}Capability~Needs:N_{c_{j}}=\sum_{i=1}^{c_{j}}C_{i}\cdot\mathbb{P}(C_{i}|X_{j},T,N_{b_{j}});\end{split}$ (6)
$\begin{split}Teaming~~Needs:N_{t_{j}}=\sum_{i=1}^{t_{j}}T_{i}\cdot\mathbb{P(}T_{i}|X_{j},T,N_{c_{j}});\end{split}$ (7)
$\begin{split}Learning~~Needs:N_{l_{j}}=\sum_{i=1}^{l_{j}}L_{i}\cdot\mathbb{P(}L_{i}|X_{j},T,N_{t_{j}});\end{split}$ (8)
Here, $X_{j}=\{P_{j},C_{j}\}$ $\in$ $\Psi$ is the combined state of the agent $j$ with $P_{j}$ being the perceived information by that agent and $C_{j}$ representing the communicated data from other agents. $T$ is the assigned task (goal or objective). $S_{i}$, $B_{i}$, $C_{i}$, $T_{i}$, and $L_{i}$ are the utility values of corresponding feature/factor $i$ in the corresponding levels - Safety, Basic, Capability, Teaming, and Learning, respectively. $s_{j}$, $b_{j}$, $c_{j}$, $t_{j}$, and $l_{j}$ are the sizes of agent $j$’s feature space on the respective levels of needs.
The collective need of an agent $j$ is expressed as the union of needs at all the levels in the needs hierarchy as in Eq. (9). Where each level of category needs is combined with various similar needs (expected values), presenting as a set, consisting of an individual hierarchical and compound needs matrix $N_{j}$. Fig. 12 illustrates a human-like motivation mechanism in a robot, represented as a behavior tree (BT).
$N_{j}=N_{s_{j}}\cup N_{b_{j}}\cup N_{c_{j}}\cup N_{t_{j}}\cup N_{l_{j}}$ (9)
Where the needs represent motivations or goals of agents to achieve various tasks and consist of their value systems in the development. Furthermore, when AI agents work with humans, they should satisfy humans’ needs and assist humans in fulfilling their missions efficiently. Especially trust among agents can be evaluated based on their current status and needs, performance, knowledge, experience, motivations, etc.
According to the above discussion, in the following sections, we will review the existing literature from the perspectives of single-agent systems, multi-agent systems (MAS), trust among agents, and human-robot interaction (HRI) from the utility-oriented cognitive modeling angle.
3 Single-Agent Systems